The length of the string of a simple pendulum is measured with a metre scale to be 90.0 cm. The radius of the bob plus the length of the hook is calculated to be 2.13 cm using measurement with a slide calipers. What is the effective length of the pendulum ? (The effective length is defined as the distance between the point of suspension and the centre of the bob.)

Grade:11

1 Answers

Kevin Nash
9 years ago
As shown in the figure, Actual effective length = (90.0 + 2.13) cm But, in the measurement 90.0 cm, the no. of significant digits is only 2. So, the addition must be done by considering only 2 significant digits of each measurement. So, effective length = 90.0 + 2.1 = 92.1 cm.